OpenStudy (anonymous):

If f(x) = x + 5, what does f(8) represent? The value if x when (x + 5) = 8 The value of 8f(x) when x = 8 The value of f(x) when (x + 5) = 8 The value of (x + 5) when x = 8

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

im thinking its The value of f(x) when (x + 5) = 8

jigglypuff314 (jigglypuff314):

not quite :) The value of f(x) when (x + 5) = 8 would be f(x) = 8 not f(8) f(8) means when x = 8

OpenStudy (anonymous):

so it would be B, im confuzed

OpenStudy (anonymous):

i get it now but i still dont know if im getting the right answer (The value of (x + 5) when x = 8)?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that helps alot!!